OpenStudy (anonymous):

Michael has a drawer with 8 pairs of black socks and 12 pairs of white socks. Without looking he takes a white pair of socks out of the drawer. What is the probability that the next pair he takes out is black?

OpenStudy (anonymous):

8/20=.4 .4= 40%

OpenStudy (anonymous):

Reasoning to my answer: 8 because there are 8 pairs of black socks 20 because 8+12=20 (you have to add them up to get a full 100%) then i divided it and got 40% hope that helped!

OpenStudy (anonymous):

Thanks it did !

OpenStudy (anonymous):

welcome!